# Dynamics of Pseudoentanglement

### v1

@@ V Generation of Pseudoentanglement
Figure 9: Local automaton circuit model for the generation of pseudoentanglement from product states. (a) Circuit model: the same as Fig. 5(a) but with a product initial state |$\ket{0}^{\otimes N-N_{A}}\otimes\ket{+}^{\otimes N_{A}}$ (no random phases). (b) Schematic summary of bounds for the pseudothermalization time. Solid lines indicate unconditional bounds while the dot-dashed line depends on Conjecture 1. Results are the same as in Fig. 5(b) but with a modified upper bound $\tilde{O}(m^{2}N^{2})$, Eq. (44). (c) Given the more complex mapping between quantum ensemble distinguishability and the mixing of classical Markov chains, Eq. (38-39), the precise nature of pseudothermalization (e.g., whether there is a cut-off) remains unclear in this case.
@@
Thus far we have discussed the propagation of pseudoentanglement: we have taken initial states that already contain random subset-phase ensembles on subsystems, and studied how these turn into random subset-phase states on the whole system under automaton circuit dynamics. This approach was motivated partly by analogy with the problem of entanglement spreading in quantum quenches, and partly by technical simplifications that arise when random phases are already provided by the input states. In this section, we address the question of generating pseudoentanglement from initial product states. To avoid complications associated with random phases, we choose to focus on the random subset states introduced in Ref. Giurgica-Tiron and Bouland (2023) and reviewed in Sec. II.1. We use only the local gate set of Sec. IV.3.
@@
|$\ket{\psi_{S_{0}}}=\ket{0}^{\otimes N-k}\otimes\ket{+}^{\otimes k},$ (37)
corresponding to subset $S_{0}=\{\mathbf{z}:\ z_{1}=\dots=z_{N-k}=0\}$ of cardinality $K=2^{k}$, shown by Fig. 9(a). To be pseudoentangled, random subset ensembles need to have $\omega({\rm poly}(N))\leq K\leq o(2^{N}/{\rm poly}(N))$, therefore we take the number of qubits initially in superposition, $k$, to be $\omega(\log N)\leq k<cN$ with $10<c<$ 1 a constant. After $t$ time steps, we have a subset state |$\ket{\psi_{S_{t}}}$ where $S_{t}$ is a random variable on $\Sigma_{K}$ distributed according to a time-evolved distribution $p_{t}$. To characterize the generation of pseudoentanglement, we look at the trace distance between $m$-th moment operators, $\|\rho_{t}^{(m)}-\rho_{{\rm Haar}}^{(m)}\|_{{\rm tr}}$, with $\rho_{t}^{(m)}=\sum_{S}p_{t}(S)\ket{\psi_{S}}\bra{\psi_{S}}^{\otimes m}$. As long as $m\leq O({\rm poly}(N))$, the application of Proposition 2 yields
$\|\rho_{t}^{(m)}-\rho_{{\rm Haar}}^{(m)}\|_{\rm tr}$ $\simeq\left\|\mathcal{M}^{(m)}_{p_{t}}\right\|_{\rm tr},$ (38)
@@
is the deviation from uniformity of a Markov chain on $\Sigma_{m+\delta}$ induced by the random automaton circuit, as in Sec. IV.
In this case, the pseudothermalization problem is not directly mapped onto the mixing of a single classical random walk. Rather, it is mapped in a nontrivial way, via Eq. (39), to the mixing of $m+1$ distinct random walks, indexed by $\delta=0,1,\dots m$, and distributed according to $\Phi_{m+\delta\leftarrow K}[p_{t}]$, defined by Eq. (16).
More specifically, the diagonal terms in $\mathcal{M}^{(m)}_{p_{t}}$ correspond to $\delta=0$, i.e., to the walk on $\Sigma_{m}$ distributed as $\Phi_{m\leftarrow K}[p_{t}]$. This is the same walk that controlled the spreading of pseudoentanglement in Sec. IV. However, the off-diagonal terms, corresponding to $\delta>0$, make this problem more complex. While we do not have an exact solution in this case, one can still obtain nontrivial bounds on the pseudothermalization time $t_{\rm PTherm}^{(m)}$, as we discuss next.
First, a lower bound on $t_{\rm PTherm}^{(m)}$ is obtained by using the monotonicity of trace norm under quantum channels: $\|\mathcal{D}(\mathcal{M})\|_{\rm tr}\leq\|\mathcal{M}\|_{\rm tr}$, where $\mathcal{D}$ is a dephasing super-operator that annihilates off-diagonal matrix elements: $[\mathcal{D}(\mathcal{M})]_{S^{\prime},S^{\prime\prime}}=\delta_{S^{\prime},S^{\prime\prime}}\mathcal{M}_{S^{\prime},S^{\prime}}$. This gives
$\|\rho_{t}^{(m)}-\rho_{{\rm Haar}}^{(m)}\|_{\rm tr}\geq 2\|\Phi_{m\leftarrow K}[p_{t}]-\pi_{m}\|_{\rm tv}$ (41)
@@
$t_{\rm PTherm}^{(m)}\geq t_{\rm mix}^{(m)}.$ (42)
Here $t_{\rm mix}^{(m)}$ is the mixing time of the Markov chain $\Phi_{m\leftarrow K}[p_{t}]$, studied in Sec. IV; the lower bounds we derived there thus apply directly to the pseudothermalization time in the present problem (in one spatial dimension):
$t_{\rm PTherm}^{(m)}\geq\Omega(N^{2},mN/\log(N)).$ (43)
Furthermore, since the matrix $\mathcal{M}^{(m)}_{p_{t}}$ that controls pseudothermalization, Eq. (39), consists of differences $\Phi_{m+\delta\leftarrow K}[p_{t}]-\pi_{m+\delta}$, we can guarantee pseudothermalization once all the walks ($\delta=0,1,\dots m$) have come sufficiently close to equilibrium. By properly taking into account the multiplicity of each $\delta$, we derive the following upper bound on the $m$-th pseudothermalization time (see Appendix G):
$t_{\rm PTherm}^{(m)}\leq\tilde{O}(m^{2}N^{2}),$ (44)
where $\tilde{O}$ indicates an upper bound up to logarithmic factors. Again, this is conditional on Conjecture 1.
@@
    at early times $t\leq\tilde{O}(N^{2},mN)$, the system can be efficiently distinguished from Haar-random at the $m$-copy level.
@@
VI Discussion
VI.1 Summary
In this work, we discussed the dynamical generation and propagation of pseudoentanglement. We considered automaton circuits as models of dynamics that preserve the relevant ensembles of states: both the subset-phase states and the subset states, which have been introduced as “standard models” of pseudoentanglement in previous work Aaronson et al. (2023); Giurgica-Tiron and Bouland (2023). Specifically, we considered two models of automaton dynamics: an all-to-all interacting circuit, and a circuit of geometrically-local, 3-body, Toffoli-like gates on a $d$-dimensional qubit lattice (focusing mostly on $d=1$). The all-to-all interacting model is exactly solvable, but the time scales for generation and propagation of pseudoentanglement are unphysically large (exponential in system size); the local circuit model does not suffer from these exponential time scales, but is less analytically tractable, so that our results are limited to a combination of rigorous bounds and conjectures supported by numerics.
@@ VI.1 Summary
We separately analyzed two related problems: the spreading of pseudoentanglement when a small pseudoentangled subsystem is placed in contact with a large trivial system, and the generation of pseudoentanglement starting from a product state. In both cases the key technical step involves mapping the problem onto the equilibration of certain classical random walks over subsets of the computational basis. This creates a connection to an interesting feature of Markov chains on highly-connected spaces: the cutoff phenomenon, whereby mixing happens in a step-like fashion in the limit of large system size, giving asymptotically sharp transitions to equilibrium. We argued that our Markov chains in many cases exhibit a cutoff phenomenon, giving sharp dynamical transitions in the distinguishability of our output states from Haar-random ones.
We named the dynamical approach toward pseudoentangled ensembles pseudothermalization, motivated by the analogy with the role of entanglement in quantum thermalization. This new concept adds to a growing body of research on the nuances of thermalization, chaos, ergodicity, scrambling, etc, that can be revealed only by higher-moment statistical quantities, thus going beyond the first- or second-moment quantities (e.g., local expectation values Srednicki (1994), out-of-time-ordered correlators von Keyserlingk et al. (2018)) that are traditionally the basis of such notions.
@@
Other recently introduced ideas in this direction include deep thermalization Cotler et al. (2023); Ho and Choi (2022); Ippoliti and Ho (2023, 2022); Lucas et al. (2023); Bhore et al. (2023); Shrotriya and Ho (2023); Chan and De Luca (2024), complete Hilbert-space ergodicity Pilatowsky-Cameo et al. (2023, 2024), and thermalization via free probability or cumulants Fava et al. (2023); Pappalardi et al. (2023). However, those are stronger notions, based on information-theoretic (pseudo-)randomness of certain state ensembles generated for example by quantum measurements or by quasiperiodic drive sequences. The randomness can be quantified through the notion $\epsilon$-approximate quantum state $m$-design, corresponding to the trace distance between moment operators being smaller than a tolerance $\epsilon$ (as in Eq. (2)); in those stronger versions of thermalization, one generally requires or expects an exponentially small tolerance $\epsilon\sim 1/\exp(N)$, or possibly even exact state designs ($\epsilon=0$). For $m\geq 2$, these conditions imply high entanglement. Pseudothermalization in contrast is based on a computational notion of (pseudo-)randomness, where the observer is limited to polynomial resources (e.g. number of state copies) in their attempt to tell the states apart from truly random ones. Again in the language of state designs, this corresponds to a tolerance $\epsilon=o(1/{\rm poly}(N))$ that is superpolynomially small (in order to trick the polynomially-bound observer), but not exponentially small, thus allowing for weakly entangled states.
VI.2 Outlook

### v3

@@ 5 Generation of Pseudoentanglement
Figure 9: Local automaton circuit model for the generation of pseudoentanglement from product states. (a) Circuit model: the same as Fig. 5(a) but with a product initial state |$\ket{0}^{\otimes N-N_{A}}\otimes\ket{+}^{\otimes N_{A}}$ (no random phases). (b) Schematic summary of bounds for the $m$-copy indistinguishability time. Solid lines indicate unconditional bounds while the dot-dashed line depends on Conjecture 1. Results are the same as in Fig. 5(b) but with a modified upper bound $\tilde{O}(m^{2}N^{2})$, Eq. (45). (c) Given the more complex mapping between quantum ensemble distinguishability and the mixing of classical Markov chains, Eq. (39-40), the precise nature of ensemble pseudothermalization (e.g., whether there is a cut-off) remains unclear in this case.
@@
Thus far we have discussed the propagation of pseudoentanglement: we have taken initial states that already contain random subset-phase ensembles on subsystems, and studied how these turn into random subset-phase states on the whole system under automaton circuit dynamics. This approach was motivated partly by analogy with the problem of entanglement spreading in quantum quenches, and partly by technical simplifications that arise when random phases are already provided by the input states. In this section, we address the question of generating pseudoentanglement from initial product states. We use only the local gate set of Sec. 4.3. To avoid technical complications associated with random phases, we choose to focus on the random subset states introduced in Refs. giurgica-tiron_pseudorandomness_2023 ; jeronimo_pseudorandom_2024 and reviewed in Sec. 2.1. We remark that this is a choice based on technical convenience. In principle an automaton circuit interspersed by phase gates will be able to produce random subset-phase states when acting on a trivial subset input states; however, we lack a statement like Propositions 1 (uniformly random phase functions) and 2 (constant phase functions) for the generic distribution of phase functions that would arise at intermediate times during that type of dynamics. However we note that the dynamics of subsets $S$ is independent of the phase functions $f$ in these automaton models, so convergence to random subset-phase states will be no faster than convergence to random subset states in the absence of phases. Thus it is reasonable to focus on random subset states in the following, while leaving the exploration of subset-phase state dynamics to future work.
@@
|$\ket{\psi_{S_{0}}}=\ket{0}^{\otimes N-N_{A}}\otimes\ket{+}^{\otimes N_{A}},$ (38)
corresponding to subset $S_{0}=\{\mathbf{z}:\ z_{1}=\dots=z_{N-N_{A}}=0\}$ of cardinality $K=2^{N_{A}}$, shown by Fig. 9(a). To be pseudoentangled according to our definition (Sec. 2.1), random subset ensembles need to have $\omega({\rm poly}(N))\leq K\leq 2^{o(N)}$, therefore we take the number of qubits initially in superposition, $k$, to be $\omega(\log N)\leq N_{A}<o(N)$. After $t$ time steps, we have a subset state |$\ket{\psi_{S_{t}}}$ where $S_{t}$ is a random variable on $\Sigma_{K}$ distributed according to a time-evolved distribution $p_{t}$. To characterize the generation of pseudoentanglement, we look at the $m$-copy distinguishability from the Haar ensemble, i.e. the trace distance $D_{\rm tr}(\rho_{t}^{(m)},\rho_{{\rm Haar}}^{(m)})$, with $\rho_{t}^{(m)}=\sum_{S}p_{t}(S)\ket{\psi_{S}}\bra{\psi_{S}}^{\otimes m}$. As long as $m\leq O({\rm poly}(N))$, the application of Proposition 2 yields
$D_{\rm tr}\left(\rho_{t}^{(m)},\rho_{{\rm Haar}}^{(m)}\right)$ $\simeq\frac{1}{2}\left\|\mathcal{M}^{(m)}_{p_{t}}\right\|_{\rm tr},$ (39)
@@
is the deviation from uniformity of a Markov chain on $\Sigma_{m+\delta}$ induced by the random automaton circuit, as in Sec. 4.
In this case, the ensemble pseudothermalization problem is not directly mapped onto the mixing of a single classical random walk. Rather, it is mapped in a nontrivial way, via Eq. (40), to the mixing of $m+1$ distinct random walks, indexed by $\delta=0,1,\dots m$, and distributed according to $\Phi_{m+\delta\leftarrow K}[p_{t}]$, defined by Eq. (17).
More specifically, the diagonal terms in $\mathcal{M}^{(m)}_{p_{t}}$ correspond to $\delta=0$, i.e., to the walk on $\Sigma_{m}$ distributed as $\Phi_{m\leftarrow K}[p_{t}]$. This is the same walk that controlled the spreading of pseudoentanglement in Sec. 4. However, the off-diagonal terms, corresponding to $\delta>0$, make this problem more complex. While we do not have an exact solution in this case, one can still obtain nontrivial bounds on the indistinguishability time $t_{\rm indist}^{(m)}$, as we discuss next.
First, a lower bound on $t_{\rm indist}^{(m)}$ is obtained by using the monotonicity of trace norm under quantum channels: $\|\mathcal{D}(\mathcal{M})\|_{\rm tr}\leq\|\mathcal{M}\|_{\rm tr}$, where $\mathcal{D}$ is a dephasing super-operator that annihilates off-diagonal matrix elements: $[\mathcal{D}(\mathcal{M})]_{S^{\prime},S^{\prime\prime}}=\delta_{S^{\prime},S^{\prime\prime}}\mathcal{M}_{S^{\prime},S^{\prime}}$. This gives
$D_{\rm tr}\left(\rho_{t}^{(m)},\rho_{{\rm Haar}}^{(m)}\right)\geq D_{\rm tv}\left(\Phi_{m\leftarrow K}[p_{t}],\pi_{m}\right)$ (42)
@@
$t_{\rm indist}^{(m)}\geq t_{\rm mix}^{(m)}.$ (43)
Here $t_{\rm mix}^{(m)}$ is the mixing time of the Markov chain $\Phi_{m\leftarrow K}[p_{t}]$, studied in Sec. 4; the lower bounds we derived there thus apply directly to the $m$-copy indistinguishability time in the present problem (in one spatial dimension):
$t_{\rm indist}^{(m)}\geq\Omega(N^{2},mN/\log(N)).$ (44)
Furthermore, since the matrix $\mathcal{M}^{(m)}_{p_{t}}$ that controls distinguishability, Eq. (40), consists of differences $\Phi_{m+\delta\leftarrow K}[p_{t}]-\pi_{m+\delta}$, we can guarantee ensemble pseudothermalization once all the walks ($\delta=0,1,\dots m$) have come sufficiently close to equilibrium. By properly taking into account the multiplicity of each $\delta$, we derive the following upper bound on the $m$-copy indistinguishability time (see Appendix G):
$t_{\rm indist}^{(m)}\leq O(\log(K)m^{2}N^{2})\leq\tilde{O}(m^{2}N^{2}),$ (45)
where the latter inequality holds if we take $K$ to be quasipolynomial, i.e., $\log K={\rm polylog}(N)$. Again, this is conditional on Conjecture 1.
@@
    at early times $t\leq\tilde{O}({\rm max}(N^{2},mN))$, the system can be efficiently distinguished from Haar-random at the $m$-copy level.
@@
6 Discussion
6.1 Summary
In this work we have examined the role of entanglement in the thermalization of isolated quantum many-body systems from a new perspective. We have introduced models of quantum dynamics that at late times reproduce all efficiently observable predictions of infinite-temperature quantum equilibrium, while only generating a relatively small amount of entanglement. This is contrary to the expectation that high-temperature equilibrium should be described by nearly-maximally-entangled random states, following the general “maximum entropy principle” that underpins thermodynamics. This counterintuitive result is made possible by the existence of pseudoentanglement, i.e., the ability of certain quantum state ensembles to securely hide their true entanglement structure from observers with limited resources.
We have constructed random circuit models that provably equilibrate to ensembles of pseudoentangled states at late times, a phenomenon we have named “ensemble pseudothermalization”. To understand this process quantitatively, we have studied the dynamical generation and propagation of pseudoentanglement in these circuit models through a mix of exact results, rigorous bounds, and numerical simulations. More specifically, our models are based on random reversible classical circuits that preserve certain types of states (subset-phase states aaronson_quantum_2023 and subset states giurgica-tiron_pseudorandomness_2023 ; jeronimo_pseudorandom_2024 ) known to form pseudoentangled ensembles. We separately analyzed two related problems: the spreading of pseudoentanglement when a small pseudoentangled subsystem is placed in contact with a large trivial system, and the generation of pseudoentanglement starting from a product state.
@@ 6.1 Summary
The key technical step in our analysis involves mapping the distinguishability between quantum state ensembles onto the equilibration of certain classical random walks over subsets of the computational basis. This creates a connection to an interesting feature of Markov chains on finite sets: the cutoff phenomenon, whereby mixing happens in a step-like fashion in the limit of large system size, giving asymptotically sharp transitions to equilibrium. We argued that our Markov chains in many cases exhibit a cutoff phenomenon, giving sharp dynamical transitions in the distinguishability of our output states from Haar-random ones.
@@
The new concept of ensemble pseudothermalization adds to a growing body of research on the nuances of thermalization, chaos, ergodicity, scrambling, etc., that can be revealed only by higher-moment statistical quantities, thus going beyond the first- or second-moment quantities (e.g., local expectation values srednicki_chaos_1994 , out-of-time-ordered correlators von_keyserlingk_operator_2018 ) that are traditionally the basis of such notions. Other recently introduced ideas in this direction include deep thermalization cotler_emergent_2023 ; ho_exact_2022 ; ippoliti_dynamical_2023 ; ippoliti_solvable_2022 ; lucas_generalized_2023 ; bhore_deep_2023 ; shrotriya_nonlocality_2023 ; chan_projected_2024 , complete Hilbert-space ergodicity pilatowsky-cameo_complete_2023 ; pilatowsky-cameo_hilbert-space_2024 , and thermalization via free probability or cumulants fava_designs_2023 ; pappalardi_general_2023 . Those notions tend to be stronger than standard thermalization (as defined from the equilibration of local reduced density matrices). They are based on information-theoretic pseudorandomness of certain state ensembles generated for example by quantum measurements or by quasiperiodic drive sequences. The randomness can be quantified through the notion $\epsilon$-approximate quantum state $m$-design, corresponding to the trace distance between $m$-th moment operators being smaller than a tolerance $\epsilon$ (as in Eq. (3)); in those stronger versions of thermalization, one generally requires or expects an exponentially small tolerance $\epsilon\sim 1/\exp(N)$, or possibly even exact state designs ($\epsilon=0$). For $m\geq 2$, these conditions imply high entanglement.
Ensemble pseudothermalization in contrast is based on a computational notion of pseudorandomness, where the observer is limited to polynomial resources (e.g. time, number of state copies) in their attempt to tell the states apart from truly random ones. Again in the language of state designs, this corresponds to a tolerance $\epsilon=o(1/{\rm poly}(N))$ that is superpolynomially small (in order to trick the polynomially-bound observer), but not exponentially small, thus allowing for weakly entangled states. Moreover, while we have used the trace distance as a measure of distinguishability in this work, in reality the observer might be limited to sub-optimal measurements that do not saturate the trace distance bound (the optimal distinguishing measurement might require a deep quantum circuit to implement). It would be interesting to explore how our results would change when keeping this additional limitation into account.
6.2 Implications
Our results carry several implications for quantum physics and quantum information science. We summarize them below, while noting that fully unraveling their consequences will require future work.
(i) Learning quantum processes and entanglement growth. Ensemble pseudothermalization implies limits on the ability of an observer to efficiently characterize entanglement growth in many-body dynamics, an important tool in experimental studies of thermalization kaufman_quantum_2016 . While pseudoentanglement makes this statement for individual states, our work translates this statement to time evolution: it is possible to efficiently verify, e.g., the linear growth of quantum entanglement versus time $t$ (a feature of quantum chaotic dynamics) up to a point, $t\sim\log N$, but not beyond it. Our results also have broader implications on the learning of quantum processes. While our automaton circuits are not pseudorandom (evident from their trivial action on computational basis input states), they would nonetheless appear random to several protocols for quantum process learning huang_learning_2023 ; levy_classical_2024 . These protocols are based on feeding in random Pauli product states (randomly ±$\pm 1$ in the $X$, $Y$ or $Z$ basis on each qubit) and then performing randomized Pauli-basis measurements on the output. Since random Pauli product states are subset-phase states of typical cardinality $K\sim 2^{(2/3)N}$ (each $X$ or $Y$ basis state doubles the cardinality of the subset), this ensemble of input states can pseudothermalize for any polynomial number of copies $m$. Classical shadows on the output states will not be able to tell the state apart from Haar states, and thus the process cannot be efficiently distinguished from a Haar-random unitary by these methods.
(ii) Detection of measurement-induced phenomena. Much research has been devoted lately to phenomena that take place in conditional, post-measurement states of quantum many-body systems. These include measurement-induced entanglement and teleportation transitions bao_finite-time_2024 ; hoke_measurement-induced_2023 as well as ‘deep thermalization’, the emergence of universal wavefunction distributions on subsystems from measurements on the rest of the system (the ‘projected ensemble’) cotler_emergent_2023 ; ho_exact_2022 . Given the intrinsically non-deterministic nature of measurement, detection of all these phenomena naïvely relies on postselection, which is inefficient for many-body states; however, efficient detection may become possible when allowing adaptive and computationally-enhanced strategies gullans_scalable_2020 ; noel_measurement-induced_2022 ; hoke_measurement-induced_2023 ; garratt_probing_2024 ; mcginley_postselection-free_2024 . Ensemble pseudothermalization presents universal limits on these strategies. The subset-phase state ensembles obtained as fixed points in our dynamics lead to very simple post-measurement states: when a finite fraction of the system is measured in the computational basis, one almost always obtains a product state on the remainder of the system191919 Given a subset state |$\ket{\psi_{S}}$, measuring a bitstring $\zeta$ on subsystem $\bar{A}$ collapses the state to a superposition of all computational basis states |$\ket{\eta}_{A}$ on $A$ such that the combined bitstring $z=(\eta,\zeta)$ belongs to the original subset $S$; but this typically specifies a unique $\eta$. The probability of two distinct $\eta,\eta^{\prime}$ appearing in the post-measurement state is the probabilty that two bitstrings $z,z^{\prime}\in S$ agree on the entire subsystem $\bar{A}$, which is $O(K^{2}2^{-|\bar{A}|})$ (by birthday asymptotics) for a random subset state. As long as $K=2^{o(N)}$ and $|\bar{A}|=\Theta(N)$ this probability is exponentially small in $N$. . This means, e.g., that our pseudoentangled states are in the trivial phase for measurement-induced teleportation, and have a trivial projected ensemble. One the contrary, Haar-random states are in the non-trivial (teleporting) phase bao_finite-time_2024 and their projected ensemble can form higher state designs cotler_emergent_2023 . Thus any method to efficiently diagnose these facts (without prior knowledge of the state) would serve as an efficient protocol to distinguish pseudoentangled states from Haar-random states, which is by definition impossible. It follows that any efficient protocol for the detection of these measurement-induced phenomena must use prior information about the state preparation mcginley_postselection-free_2024 . Without this information, no efficient protocol (even allowing for arbitrary adaptivity and multi-copy measurements) can succeed.
(iii) Fundamental physics implications. Our results highlight the fact that extensive amounts of entanglement are generally not necessary to reproduce the predictions of high-temperature thermal equilibrium. The results on the generation of pseudoentanglement from product states are especially remarkable: they show that a disentangled product state, featuring only a modest amount (vanishing density) of coherence and acted upon by ‘classical’ dynamics and [see Fig. 9(a)], becomes indistinguishable from a near-maximally entangled random state. This raises the prospect that the universe might, in reality, be “less quantum” than we think—for example governed by mostly classical local update rules, enhanced by relatively small amounts of quantum resources like entanglement and coherence. Theory proposals of “emergent quantum mechanics” slagle_testing_2022 may include a cut-off on the allowed amount of entanglement in many-body systems, or present deviations from standard quantum mechanics only in highly entangled states. Our findings suggest that it may be impossible to efficiently rule out such scenarios in general.
6.3 Outlook